OpenStudy (anonymous):

"Line A and Line B are parallel. The slope of Line A is -2. Find the value of x if the slope of Line B is 3x+3." Why doesn't that make sense?

OpenStudy (anonymous):

Yea, my teacher gave me this question and it had me stumped for a while. Thanks!

Directrix (directrix):

Did the "why doesn't this make sense" come with the problem? If not, then I think there is a better answer to the question.

Directrix (directrix):

slope of Line B is 3x+3 (3x +3) = -2 because A and B are parallel.

OpenStudy (anonymous):

I wrote that part. Good to know I wasn't wrong.

Directrix (directrix):

You were wrong, I think.

Directrix (directrix):

(3x +3) = -2 because A and B are parallel. 3x = -2 - 3 3x = -5 x = -5/3 So, x = -5/3 is the value of x that makes the blope of line B = -2 which means B and line A are parallel.

OpenStudy (anonymous):

Oh. I get it. Sorta. Thanks for the help.

Directrix (directrix):

Question?

OpenStudy (anonymous):

Don't worry, I think I got it now.